A machine that manufactures automobile parts is estimated to produce defective parts 6% of the time. If this estimate is correct, and 6 parts produced by this machine are randomly selected, what is the probability that fewer than 2 turn out to be defective? Round your answer to four decimal places.

Answer by stanbon(75887)

A machine that manufactures automobile parts is estimated to produce defective parts 6% of the time. If this estimate is correct, and 6 parts produced by this machine are randomly selected, what is the probability that fewer than 2 turn out to be defective? Round your answer to four decimal places.
------------------
Binomial Problem with n = 6 and p(defect)=0.06 and p(not defect)= 0.94
------------------
P(0<= x <=1) = P(x=0 defective) + P(x=1 defective)
---
= 6C0(0.06)^0*(0.94)^6 + 6C1(0.06)*(0.94)^5
----
= (0.94)^6 + 6*0.06*(0.94)^5
---
= 0.6899 + 0.2642
----------
= 0.9541
===============
Cheers,
Stan H.
===============